to find the degeneracy of excited state of nitrogen atom
Answer #1
The degeneracy of an excited state of nitrogen atom having electronic configuration
1s22s22p23d1
For the 3d1 state
S = (1 / 2) for the single electron
L = 2 for the d - orbital
Therefore, the degeneracy of the state is given by
= (2L+1)(2S+1) = (2*2 +1)(2* (1 / 2)+1)
= (5)(2)
= 10
answered by: Carl
Answer #2

Configuration : 1s^22s^22p^23d1


Degeneracy of a shell (g) = Y! /  X!(Y - X)!


Where X is no. of electrons and Y is total no. of electron occupy in that shell.


For 1s X = 2, Y = 2 ; g = 1

For 2s X = 2, Y = 2 ; g = 1

For 2p X = 2, Y = 6 ; g = 15

For 3d X=1,   Y=10;   g=10

So, G = 1*1*15*10= 150